OpenStudy (anonymous):

2+2=5 proof My teacher said we must prove 2+2=5 there is a way and he wants us to find out.

OpenStudy (calculusxy):

x = y x^2 = xy x^2-y^2 = xy-y^2 (x+y)(x-y) = y(x-y) x+y = y 2y = y 2 = 1 1 = 0 Now since 2+2 = 2+2+0 and we apply the lemma proof so that 0 = 1, 2+2+0 = 2+2+1 =5 THEREFORE 2+2 = 5

OpenStudy (calculusxy):

That is called the Robin Lemma proof, I believe.

OpenStudy (calculusxy):

@Futuristicmario

OpenStudy (tkhunny):

There is not a way that is without error, but it shouldn't be too hard. 0 = 1 is pretty common. Adding 4 to that might easily produce 4 = 5. The usual might be: 1) Start with two unequal numbers, a and b 2) c = (a+b)/2 - the average of a and b 3) The 2c = a+b 4) Since a-b is not zero 2c(a-b) = (a+b)(a-b) ==> 2ac - 2bc = a^2 - b^2 5) Algebra: a^2 - 2ac = b^2 - 2bc 6) Add c^2: a^2 - 2ac + c^2 = b^2 - 2bc + c^2 7) Factor: (a-c)^2 = (b-c)^2 8) Square Root: a-c = b-c 9) Add c: a = b 10) Take a = 5 and b = 4, which are NOT equal, for 1), and we have from 9) 5 = 4 = 2+2. Wasn't that fun!? Please Note: There ARE errors in there.

OpenStudy (anonymous):

@calculusxy :??? (x+y)(x-y) = y(x-y) ???? (x+y)(x-y)=x^2-y^2 I think !:)

OpenStudy (unklerhaukus):

for large values of two, eg 2.3 2.3+2.3 =4.6 = 5 , a small value for five